OpenStudy (brittany16):

Solve for the variable. b + (-14) = 25

OpenStudy (anonymous):

add 14 to both sides b=39

jimthompson5910 (jim_thompson5910):

b + (-14) = 25 b -14 = 25 b -14 +14 = 25 + 14 b = 39